# Active Causal Structure Learning with Latent Variables: Towards Learning to Detour in Autonomous Robots

### v1

@@ Appendix 0.A Divergence surprise proofs
$D_{KL}(Q||P)\xrightarrow{}0$Leftrightarrow $D_{S}(Q||P)^{2}\xrightarrow{}0$.
@@
$=\frac{D_{KL}(Q||P)^{2}}{V_{I}(P)}+\frac{2D_{KL}(Q||P)}{V_{I}(P)}(H(Q)-H(P))+\frac{(H(Q)-H(P))^{2}}{V_{I}(P)}=$
$=A\;+\;B\;+\;C.$
@@
[…]
@@
[…]
@@
$=\frac{1}{n-1}\sum_{i=1}^{n}\left[-log(P(X=X_{i}))-H(P)+H(P)-\frac{1}{n}\sum_{j=1}^{n}-log(P(X=X_{j}))\right]^{2}=$
$=\frac{1}{n-1}\sum_{i=1}^{n}\left[-log(P(X=X_{i}))-H(P)-\frac{1}{n}\left(\sum_{j=1}^{n}-log(P(X=X_{j})-H(P)\right)\right]^{2}.$
@@
$=\frac{n}{n-1}\left[\frac{1}{n}\sum_{i=1}^{n}D_{i}^{2}-\bar{D}^{2}\right].$
@@
[…]
@@
[…]
@@ 0.B.1.1 Actions Space
Table 4: Correspondence between category and interval of $Step\;Forward$ discretization
@@
Table 5: Correspondence between category and interval of $Step\;Aside$ discretization
@@
To go from discrete to continuous values it has been decided to select a random decision value using a uniform distribution with the appropriate bounds, depending on the discrete decision selected by the agent. For instance, if the agent decide to do $Step\;Forward$ with a value $1$, then it will returned to the Environment to execute $SA(x,y,\alpha,s)$ with $s\sim U(0.5,1)$.
@@ 0.B.1.2 Observation Space
In the Observation Space, $Depth$ and $Heading\;Angle$ are the only observation variables that need a discretization.
For $Depth$, the interval $[0,d_{max}]$ has been divided into five equal parts, where $d_{max}=\sqrt{x_{max}^{2}+y_{max}^{2}}$
With regard to the $Heading\;Angle$, the interval $[-\pi,\pi]$ has been divided into nine equal parts.
@@ 0.B.2.2 BarrierImpact
Figure 12: $BarrierImpact$ restrictor
@@ 0.B.3 Transition Probabilities
This section outlines the transition probability matrices employed by the agent. Each matrix consists of rows representing the values of $Obs_{t}$ and columns representing the values of $Obs_{t+1}$. As each probability transition $Obs_{t+1}$ may depends on the value of $Step\;Forward$ and $Step\;Aside$, the matrices are expressed in terms of the value of $SF$ and $SA$.
0.B.3.1 $Depth$
@@
0.B.3.2 $Heading\;Angle$
$\begin{bmatrix}F_{1}&F_{1}^{-}&0&0&0&0&0&0&0&0&F_{4}^{+}\\ F_{1}^{+}&F_{1}&F_{1}^{-}&0&0&0&0&0&0&0&0\\ 0&F_{1}^{+}&F_{1}&F_{2}^{-}&0&0&0&0&0&0&0\\ 0&0&F_{1}^{+}&F_{1}&F_{2}^{-}&0&0&0&0&0&0\\ 0&0&0&F_{2}^{+}&F_{1}&F_{2}^{-}&0&0&0&0&0\\ 0&0&0&0&F_{2}^{+}&F_{1}&F_{3}^{-}&0&0&0&0\\ 0&0&0&0&0&F_{2}^{+}&F_{1}&F_{3}^{-}&0&0&0\\ 0&0&0&0&0&0&F_{3}^{+}&F_{1}&F_{4}^{-}&0&0\\ 0&0&0&0&0&0&0&F_{3}^{+}&F_{1}&F_{4}^{-}&0\\ 0&0&0&0&0&0&0&0&F_{4}^{+}&F_{1}&F_{4}^{-}\\ F_{1}^{-}&0&0&0&0&0&0&0&0&F_{4}^{+}&F_{1}\end{bmatrix}$
@@ 0.B.3.2 $Heading\;Angle$
0.B.3.3 $Target\;in\;Visual\;Field$
@@ 0.B.3.4 $Barrier\;Tactile$
$C(SF,SA)=\left\{\begin{array}[]{lcc}0.99&if&SA=5\\ 0.95&otherwise&\end{array}\right.$

### v2

@@ Appendix 0.A Divergence surprise proofs
$D_{KL}(Q||P)\xrightarrow{}0$ $\Leftrightarrow$ $D_{S}(Q||P)^{2}\xrightarrow{}0$.
@@
$\frac{D_{KL}(Q||P)^{2}}{V_{I}(P)}+\frac{2D_{KL}(Q||P)}{V_{I}(P)}(H(Q)-H(P))+\frac{(H(Q)-H(P))^{2}}{V_{I}(P)}=$
=A+B+C.=A\;+\;B\;+\;C.
@@
$\frac{S_{n}-n\mu_{\textbf{W}}}{\sqrt{n\sigma_{\textbf{W}}^{2}}}\xrightarrow{d}N(0,1),$
@@
$\frac{S_{n}-n\mu_{\textbf{W}}}{\sqrt{n\sigma_{\textbf{W}}^{2}}}=\frac{n}{\sqrt{n}}\frac{H(\hat{P},P)-H(P)}{\sqrt{V_{I}(P)}}=\sqrt{n}\cdot D_{S}(\hat{P}||P)\xrightarrow{d}N(0,1).$
@@
$\frac{1}{n-1}\sum_{i=1}^{n}\left[-log(P(X=X_{i}))-H(P)+H(P)-\frac{1}{n}\sum_{j=1}^{n}-log(P(X=X_{j}))\right]^{2}=$
$\frac{1}{n-1}\sum_{i=1}^{n}\left[-log(P(X=X_{i}))-H(P)-\frac{1}{n}\left(\sum_{j=1}^{n}-log(P(X=X_{j})-H(P)\right)\right]^{2}.$
@@
$\frac{n}{n-1}\left[\frac{1}{n}\sum_{i=1}^{n}D_{i}^{2}-\bar{D}^{2}\right].$
@@
$\bar{D}\xrightarrow{P}E_{P}[-log(P(X=X_{i})-H(P)]=0.$
Being a convergence in probability to a degenerate distribution, it is also held that $\bar{D}\xrightarrow{d}0$, and the Continuous Application Theorem can be applied with the function $f(x)=x^{2}$, obtaining:
$\bar{D}^{2}\xrightarrow{d}0\longleftrightarrow\bar{D}^{2}\xrightarrow{P}0.$
@@
$\sqrt{n}\cdot\frac{H(\hat{P},P)-H(P)}{\sqrt{\hat{V}_{I}(P)}}=\frac{V_{I}(P)}{\hat{V}_{I}(P)}\cdot\sqrt{n}\cdot\frac{H(\hat{P},P)-H(P)}{\sqrt{V_{I}(P)}}\xrightarrow{d}N(0,1).$
@@ 0.B.1.1 Actions Space
Table 4: Correspondence between category and interval of StepForwardStep\;Forward discretization
@@
Table 5: Correspondence between category and interval of StepAsideStep\;Aside discretization
@@
To go from discrete to continuous values it has been decided to select a random decision value using a uniform distribution with the appropriate bounds, depending on the discrete decision selected by the agent. For instance, if the agent decide to do StepForwardStep\;Forward with a value $1$, then it will returned to the Environment to execute $SA(x,y,\alpha,s)$ with $s\sim U(0.5,1)$.
@@ 0.B.1.2 Observation Space
In the Observation Space, $Depth$ and HeadingAngleHeading\;Angle are the only observation variables that need a discretization.
For $Depth$, the interval $0,d_{max}]$ has been divided into five equal parts, where $d_{max}=\sqrt{x_{max}^{2}+y_{max}^{2}}$
With regard to the HeadingAngleHeading\;Angle, the interval $\pi,\pi]$ has been divided into nine equal parts.
@@ 0.B.2.2 BarrierImpact
Figure 12: BarrierImpactBarrierImpact restrictor
@@ 0.B.3 Transition Probabilities
This section outlines the transition probability matrices employed by the agent. Each matrix consists of rows representing the values of $Obs_{t}$ and columns representing the values of $Obs_{t+1}$. As each probability transition $Obs_{t+1}$ may depends on the value of StepForwardStep\;Forward and StepAsideStep\;Aside, the matrices are expressed in terms of the value of $SF$ and $SA$.
0.B.3.1 DepthDepth
@@
0.B.3.2 HeadingAngleHeading\;Angle
[…]
@@ 0.B.3.2 HeadingAngleHeading\;Angle
0.B.3.3 TargetinVisualFieldTarget\;in\;Visual\;Field
@@ 0.B.3.4 BarrierTactileBarrier\;Tactile
[…]